OpenStudy (anonymous):

5√7 + 12√6 - 10√7 - 5√6

OpenStudy (australopithecus):

\[5(7)^{\frac{1}{2}} = 5\sqrt{7}\] add like terms

OpenStudy (anonymous):

(5-10)sqrt7+(12-5)sqrt6 -5sqrt7+7sqrt6 Ans

OpenStudy (australopithecus):

its just like dealing with a variable, \[(7)^{\frac{1}{2}}\] is just a number, \[5(7)^{\frac{1}{2}} = (7)^{\frac{1}{2}} + (7)^{\frac{1}{2}} + (7)^{\frac{1}{2}} + (7)^{\frac{1}{2}} + (7)^{\frac{1}{2}}\]

OpenStudy (australopithecus):

you cant add, \[(7)^{\frac{1}{2}} + (6)^{\frac{1}{2}}\] without generating a fraction, in mathematics you usually want to deal with exact numbers

OpenStudy (anonymous):

@Australopithecus so is the aswer 5√7 - 7√6 ?

OpenStudy (australopithecus):

Be careful with negatives https://www.wolframalpha.com/input/?i=5%E2%88%9A7+%2B+12%E2%88%9A6+-+10%E2%88%9A7+-+5%E2%88%9A6

OpenStudy (anonymous):

ohh ok i see. thnx